Timeline
Jul 26, 2016:
- 7:53 PM Ticket #1159 (26 Jul 07:31 Operation Resume) created by
 - Person in Charge: P. Condor. Experiment: EW_drifts 1) 08:30 - 19:30 2) …
 - 7:48 PM Ticket #1158 (26 Jul 19:45 Transmitter power measurements) created by
 - 07:50 LV ON 08:15 HV ON 09:00 TX Power sample (-70 dB coupling) Tx1 +6.8 …
 - 7:38 PM Ticket #1157 (26 Jul 19:32 Figures) created by
 - Figures of rti and spectras.
 - 7:25 PM Ticket #1156 (26 Jul 16:35 Spark gaps adjustment) created by
 - 10:37 Spark gaps adjustment 12:39 Spark gaps adjustment 14:48 Spark gaps …
 - 7:22 PM Ticket #1155 (26 Jul 10:10 Restart EW-Drift batch in ARES) created by
 - Restart ew_drift batch in pc Ares.
 - 7:19 PM Ticket #1154 (26 Jul 08:30 Start Experiment) created by
 - Start experiment with Transmitter all ok.
 - 7:18 PM Ticket #1153 (26 Jul 08:12 Digisonde turned off) created by
 - Digisonde is turned off
 
Jul 25, 2016:
- 12:28 PM Ticket #1152 (25 Jul 12:10 Antenna VSWR) created by
 - WUp 1.04 WDn 1.05 EUp 1.03 EDn 1.07
 - 10:10 AM Ticket #1151 (25 Jul 10:00 Log Created) created by
 - ISR School ISR Campaign (EW Drift only) Tuesday 26 July 09:00 - 19:30 LT
 
Jul 18, 2016:
- 1:21 PM Ticket #1150 (01 Jun 15:30 Transmitter power measurements) created by
 - 08:00 CROWBAR 09:00 TX Power sample (-70 dB coupling) Tx1 +1.2 dBm Tx4 …
 
Jul 16, 2016:
- 8:18 AM Ticket #1149 (16 Jul 08:03 Plots) created by
 - 16 Jul 08:03 Plots
 - 8:15 AM Ticket #1148 (16 Jul 08:03 System Off) created by
 - 16 Jul 08:03 System Off
 - 7:39 AM Ticket #1147 (16 jul 07:25 Imaging Acquisition System Off) created by
 - 16 jul 07:25 Imaging Acquisition System Off
 - 7:33 AM Ticket #1146 (16 Jul 03:50 Re-star acquisition) created by
 - 16 Jul 03:50 Re-star acquisition at PC ARES
 - 7:30 AM Ticket #1145 (16 Jul 05:44 Spark gaps adjustment) created by
 - 15 Jul 21:29 Spark gaps adjustment. 15 Jul 23:34 Spark gaps adjustment. 16 …
 - 7:24 AM Ticket #1144 (15 Jul 23:59 Plots) created by
 - Plots
 
Jul 15, 2016:
- 7:59 PM Ticket #1143 (15 Jul 19:59 Operation resume) created by
 - Person in Charge: I. Manay Experiment: Ew_Drift&Imaging 1) 08:00 - 20:00 …
 - 7:53 PM Ticket #1142 (15 Jul 19:45 Plots) created by
 - 7:47 PM Ticket #1141 (15 Jul 18:00 JARS for Imaging experiment) created by
 - Imaging experiment was started for spread-F acquisition
 - 7:46 PM Ticket #1140 (15 Jul 15:10 JARS for Hydra was stopped) created by
 - Imaging experiment was setup without transmitter to acquire hydra pass for …
 - 7:43 PM Ticket #1139 (15 Jul 19:30 Transmitter power measurement) created by
 - 09:00 TX Power sample (-70 dB coupling) Tx1 +6.8 dBm D2 - Line voltage and …
 - 7:41 PM Ticket #1138 (15 Jul 19:23 Spark gaps adjustment) created by
 - 10:15 Spark gaps adjustment. 12:30 Spark gaps adjustment. 14:55 Spark gaps …
 - 8:12 AM Ticket #1137 (15 Jul 08:00 Operation Resume) created by
 - Person in Charge: P. Condor. Experiment: EW_drifts+AMISR 1) 20:00 - 08:00 …
 - 7:56 AM Ticket #1136 (15 Jul 07:50 Figures) created by
 - Figures of rtis and spectras
 - 7:13 AM Ticket #1135 (15 Jul 06:19 Imaging Off) created by
 - Imaging OFF
 - 7:12 AM Ticket #1134 (15 Jul 03:59 Imaging On) created by
 - Imaging on because spread F is appearing
 - 7:10 AM Ticket #1133 (15 Jul 07:00 Transmitter power measurements) created by
 - 00:00 TX Power sample (-70 dB coupling) Tx1 +7.0 dBm D2 -9.2 dBm Line …
 - 7:10 AM Ticket #1132 (15 Jul 05:33 Spark gaps adjustment) created by
 - 14 07 2016 22:00 Spark gaps adjustment 15 07 2016 00:31 Spark gaps …
 - 7:05 AM Ticket #1131 (15 Jul 05:35 Noise Levels) created by
 - 14 July 2016 20:25 EW DRIFTS 1:24.39 2:24.85 3:24.78 4:24.75 IMAGING …
 - 3:07 AM Ticket #1130 (15 Jul 01:05 Imaging Off) created by
 - Imaging off
 - 3:05 AM Ticket #1129 (14 Jul 23:05 Tx off) created by
 - Problem with transmitter
 - 3:04 AM Ticket #1128 (14 Jul 20:15 Restart Pc Ares) created by
 - Restart Pc Ares because it is not working
 - 2:59 AM Ticket #1127 (14 Jul 23:59 Figures) created by
 - Figures of rti and spectras
 
Jul 14, 2016:
- 11:11 PM Ticket #1126 (14 Jul 23:00 Transmitter power measurements) created by
 - 20:00 TX Power sample (-70 dB coupling) Tx1 +6.8 dBm D2 -9.3 dBm Line …
 - 8:12 PM Ticket #1125 (14 Jul 20:00 Operation Resume) created by
 - 7:32 PM Ticket #1124 (14 Jul 19:20 Transmitter power measurements) created by
 - 13 July 15:00 LV ON 15:10 HV ON 15:18 TX Power sample (-70 dB coupling) …
 - 7:03 PM Ticket #1123 (14 Jul 18:20 Imaging atenaution) created by
 - ChA 0 Ch E 2 ChB 0 Ch F 2 ChC 0 Ch G 3 ChD 0 Ch H 0
 - 7:02 PM Ticket #1122 (14 Jul 17:00 Imaging On) created by
 - 17:00 Imaging On
 - 7:01 PM Ticket #1121 (14 Jul 16:18 EW Drift On) created by
 - 16:07 EW Drift Down, Test Driver 2 16:09 EW Drift On 16:14 EW Drift Down, …
 - 6:52 PM Ticket #1120 (14 Jul 14:46 Hydra acquisition off) created by
 - 13:00 Hydra acquisition with Imaging 14:46 Hydra acquisition stopped
 - 6:49 PM Ticket #1119 (14 Jul 13:20 Transmitter on) created by
 - 13:00 Transmitter off for Hydra acquisition 13:20 Transmitter on for Hydra
 - 6:48 PM Ticket #1118 (14 Jul 11:26 Remove Interference off) created by
 - 10:54 Turned on remove interference on the software Signal Chain 11:26 …
 - 6:46 PM Ticket #1117 (14 Jul 10:30 Diode proteccion removed) created by
 - 10:30 There was an extra set of diode protección which was removed for the …
 - 6:45 PM Ticket #1116 (14 Jul 09:17 Acquisition restarted) created by
 - 09:02 Ares PC stopped 09:17 Ares PC restarted
 - 6:38 PM Ticket #1115 (14 Jul 08:00 Operation resume) created by
 - Person in Charge: A. Belleza Experiment: Ew_Drift&Imaging 1) 20:00 - 08:00 …
 - 6:37 PM Ticket #1114 (14 Jul 06:45 Noise measure) created by
 - 13 July 23:00 ChA 28.38 ChB 28.95 ChC 29.15 ChD 29.23 14 July 01:30 …
 - 8:40 AM Ticket #1113 (14 Jul 06:40 Spark Gaps adjustment) created by
 - 13 Jul 20:40 Spark gap adjustment 23:04 Spark gap adjustment 14 Jul 01:35 …
 - 8:26 AM Ticket #1112 (14 Jul 06:00 Imaging turned off) created by
 - 06:00 Imaging turned off
 - 8:24 AM Ticket #1111 (14 Jul 05:00 Spread F finished) created by
 - 03:00 Spread F started 05:00 Spread F finished
 - 8:19 AM Ticket #1110 (14 Jul 01:15 Acquisition Restarted) created by
 - 00:20 PC Ares stopped 00:50 PC Ares restarted 01:15 Acquisition program …
 - 8:05 AM Ticket #1109 (13 Jul 21:45 Channel A EW Drift) created by
 - 21:40 Channel A EW Drift drops 1 dB 21:45 Channel A EW Drift automatically …
 
Jul 13, 2016:
- 11:47 PM Ticket #1108 (13 Jul 20:00 Operation resume) created by
 - Person in Charge: I. Manay Experiment: Ew_Drift&Imaging 1) 18:00 - 20:00 …
 - 11:43 PM Ticket #1107 (13 Jul 18:00 Experiment started) created by
 - Experiment with good data started at 6:00 PM. Vipir and AMISR systems are …
 
Jul 11, 2016:
- 10:54 AM Ticket #1106 (11 Jul 10:40 Antenna) created by
 - Wup 1.10 Wdn 1.09 Edn 1.02 Eup 1.10 Mod. Hysell 1.08 Antenna Near 1.00
 - 10:53 AM Ticket #1105 (11 Jul 10:30 Start Campaign) created by
 - Start: 11 July 2016 End: 15 July 2016 User: Fabiano Rodriguez Mode: EW …
 
Note: See TracTimeline
        for information about the timeline view.
    